# Mambular: A Sequential Model for Tabular Deep Learning

### v1

@@ Results
Generally, we can confirm the strong results from the FT-Transformer architecture as introduced by Gorishniy et al., (2021) and validated by McElfresh et al., (2024). Unsurprisingly, XGboost performs comparably well on all datasets but that is slightly outperformed by the FT-Transformer. Overall, Mambular performs best on average across all datasets among evaluated models. However, Mambular similarly struggles as the other neural methods compared to XGBoost for e.g. the Wine dataset.
Table 2: Average Rank and standard deviation of the models over all datasets. The best rank is marked in bold. MambaTab refers to the model exactly as described by Ahamed and Cheng, 2024a . MambaTabT refers to transposing the axis and thus iterating over batch size. MambaTab refers to our implementation. See the details on MambaTab in the paragraph above.
Model Avg. Rank
Mambular 2.083 $\pm$ 1.037
FT-Transformer 2.417 $\pm$ 1.256
XGBoost 3.167 $\pm$ 2.577
MambaTab 4.333 $\pm$ 1.374
ResNet 4.750 $\pm$ 1.639
TabTransformer 6.222 $\pm$ 1.618
MLP 6.500 $\pm$ 1.500
MambaTab 6.583 $\pm$ 1.801
MambaTabT 7.917 $\pm$ 1.187
The detailed results for all datasets and tasks are given below (see Table 23 and 24). The results for the original MambaTab implementation, along with a discussion of our results, can be found in Appendix C. Interestingly, for the tested datasets, all neural models perform considerably worse on the Wine dataset (see Appendix for the dataset description). Similarly, XGBoost performs worse than all neural models on the Abalone and FICO datasets. Additionally, our findings suggest that both FT-Transformer and Mambular perform very well on datasets with very few categorical features (e.g., FICO, California Housing, Abalone, CPU), although leveraging embeddings and structures originally designed for discrete data inputs.
Mambular 0.018 $\pm$ 0.000 0.452 $\pm$ 0.043 0.167 $\pm$ 0.011 0.628 $\pm$ 0.010 0.035 $\pm$ 0.005 0.132 $\pm$ 0.020 0.027 $\pm$ 0.002
Models BA $\uparrow$ AD $\uparrow$ CH $\uparrow$ FI $\uparrow$ MA $\uparrow$
Mambular 0.927 $\pm$ 0.006 0.928 $\pm$ 0.002 0.856 $\pm$ 0.004 0.795 $\pm$ 0.011 0.917 $\pm$ 0.003
3.1 Distributional Regression
To further verify the applicability of Mambular to tabular problems, we conducted a small task on distributional regression (Kneib et al.,, 2023). Distributional regression describes regression beyond the mean, i.e., the modeling of all distributional parameters. Thus, Location Scale and Shape (LSS) models can quantify the effects of covariates on not just the mean but also on any parameter of a potentially complex distribution assumed for the responses. A major advantage of these models is their ability to identify changes in all aspects of the response distribution, such as variance, skewness, and tail probabilities, enabling the model to properly disentangling aleatoric uncertainty from epistemic uncertainty. While this has been a common standard in classical statistical approaches (Stasinopoulos and Rigby,, 2008), it has not yet been widely adopted by the ML community. Recent interpretable approaches (Thielmann et al., 2024a, ), however, have demonstrated the applicability of distributional regression in tabular deep learning. Furthermore, approaches like XGBoostLSS (März,, 2019; März and Kneib,, 2022) demonstrate that tree-based models are capable of effectively solving such tasks. Below, we show that Mambular for Location Scale and Shape (MambularLSS) outperforms XGBoostLSS in terms of Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery,, 2007) when minimizing the negative log-likelihood while maintaining a small MSE. See appendix D.1 for a short introduction to the CRPS metric.
@@ 4 Ablation
We test out different architectures for Mambular to analyze whether different pooling techniques or bi-directional processing can be advantageous or may decrease performance. Sum-pooling, Max-pooling, standard Average-pooling, and Last token pooling—where only the last token in the sequence is passed to the task-specific model head—are tested. Additionally, the effectiveness of a learnable interaction layer between the features in improving performance is evaluated. This layer captures and models interactions between features by learning an interaction matrix $\mathbf{W}\in\mathbb{R}^{J\times J}$ such that $\text{interactions}=\mathbf{zW}$, where $\mathbf{z}$ is the input feature matrix, before being passed through the SSM. In Transformer networks for natural language processing, pooling via [cls] token embeddings is quite common (Gorishniy et al.,, 2021; Thielmann et al., 2024b, ). This has also been shown to be beneficial in tabular problems (Thielmann et al., 2024b, ). Instead of appending the [cls] token to the beginning of each sequence, it is appended to the end of each sequence. Pooling is then performed as last token pooling.
@@
Interestingly, we find that the vanilla architecture of average pooling, no interaction, and only one-directional processing performs consistently best among model configurations. Additionally, that last token pooling and [CLS] token pooling perform significantly worse on two of the four tested datasets. A 5-fold cross-validation is performed, using the same hyperparameters across all models The matrices $\mathbf{A},\mathbf{B}$ and $\Delta$ are not shared in bi-directional processing but each direction has its own set of learnable parameters. Thus, the bi-directional models have additional trainable parameters.
@@
Additionally, we analyze whether the order of the sequence and the position of the variables within the sequence affect the model’s performance. For textual data, shuffling the order of words/tokens has a significant impact, and even switching out single words can lead to completely different contextualized embeddings. Since these contextualized representations are pooled and fed directly to the task-specific head in Mambular, this could also affect performance. We test two different shuffling settings: i) shuffling the embeddings after they have passed through the embedding layer, ii) shuffling the sequence of variables before being passed through the embedding layers.
@@
To verify these results, we tested it on 4 of the benchmark datasets. We again find, that flipping the sequence before or after the embedding layer does not have a significant impact. Additionally, we generally find that the sequence ordering does not significantly affect model performance at the 5% level for most datasets.
[…]
Model BA $\uparrow$ AD $\uparrow$ AB $\downarrow$ CA $\downarrow$
[…]
However, for the california housing dataset we find one randomly shuffled sequence that does perform worse than the baseline at the 5% significance level. Analyzing the variable sequence, however, we do find a consistent, but puzzling pattern. The position of the variables Longitude and Latitude seems to directly influence model performance. However, Figure 4 shows that there are much stronger correlations between the features than the one between Longitude and Latitude. Additionally, we analyzed the effect strengths of all pairwise interactions with a simple Linear regression and XGBoost. We find, that the interaction effect between Longitude and Latitude is much less important than that between Longitude and Median Income for example. We have analyzed the California housing dataset in detail in Appendix D.
Figure 4: Correlation plot for California Housing dataset
This finding, while only relevant for one of the four tested real world datasets, suggests that variable ordering can have an effect on model performance for Mambular. While the default ordering demonstrates strong performance, exploring alternative orderings might further enhance Mambular’s benchmark results. As expected, we do not find these results for Attention based model. See the appendix for the results for the FT-Transformer.
Table 9: Analysis of results for CA Housing
@@
5 Limitations
The presented model is tested on a variety of datasets and benchmarked against a range of models. Note that we do not perform hyperparameter tuning, as results from Grinsztajn et al., (2022) and Gorishniy et al., (2021) indicate that most models already perform well without tuning. These studies suggest that while hyperparameter tuning does improve performance across all models simultaneously, the relative ranking of the models remains largely unchanged. This implies that a model performing best or worst with default configurations will likely maintain its ranking even after extensive tuning. Additionally, McElfresh et al., (2024) found similar results, reinforcing the idea that hyperparameter tuning benefits most models to a similar degree without altering their comparative performance.
The lack of tuning leaves room for improvement across all models. However, the default configurations for the comparison models are thoroughly tested in a variety of studies, and it is expected that if any model might benefit more from hyperparameter tuning, it would be Mambular, as we lack extensive literature to guide its default settings. For the comparison models, we based our choices on literature to ensure meaningful and high-performing parameters by default. We were able to reproduce average results from studies such as Gorishniy et al., (2021) and Grinsztajn et al., (2022). Additionally, critical hyperparameters such as learning rate, patience and number of epochs are shared among all models for a more consistent approach. All hyperparameter configurations are listed in Appendix C and can also be found at https://github.com/basf/mamba-tabular.
6 Conclusion
We present Mambular, a new architecture for tabular deep learning. We show that a truly sequential model can be applied to tabular problems, offering a fresh perspective on how to interpret and handle tabular data by treating it as a sequential problem. Our results demonstrate that a sequential model is effective for both regression and classification tasks across various datasets. The performance of Mambular, along with its extension to MambularLSS, demonstrates its broad applicability to a range of tabular tasks.
Although Mamba is still relatively new compared to architectures like the Transformer, its rapid adoption suggests significant potential for further improvement. Advances such as those proposed by Lieber et al., (2024) and Wang et al., (2024) could be particularly beneficial for tabular applications. Additionally, exploring the optimal ordering of features or incorporating column-specific information through textual embeddings could further enhance performance.
Interpreting tabular data as a sequence presents significant advantages for feature incremental learning. New features can be appended directly to the sequence without the need to retrain the entire model.
Acknowledgements
We sincerely acknowledge and appreciate the financial support provided by the Key Digital Capability (KDC) for Generative AI at BASF and the BASF Data & AI Academy, which played a critical role in facilitating this research.
Funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within project 450330162 is gratefully acknowledged.
Appendix A Datasets
Table 10: The used datasets for benchmarking. All datasets are taken from the UCI Machine Learning repository. #num and #cat represent the number of numerical and categorical features respectively. The number of features thus determines for Mambular the ”sequence length”. The train, test and val numbers represent the average number of samples in the respective split for the 5-fold cross validation. Ratio marks the percentage of the dominant class for the binary classification tasks.
Name Abbr. #cat #num train test val ratio
Regression Datasets
Diamonds DI 4 7 34522 10788 8630 -
Abalone AB 1 8 2673 835 668 -
California Housing CA 1 9 13210 4128 3302 -
Wine Quality WI 0 12 4158 1299 1039 -
Parkinsons PA 2 20 3760 1175 940 -
House Sales HS 8 19 13832 4322 3458 -
CPU small CPU 0 13 5243 1638 1310 -
Classification Datasets
Bank BA 13 8 28935 9042 7233 88.3%
Adult AD 9 6 31259 9768 7814 76.1%
Churn CH 3 9 6400 2000 1600 79.6%
FICO FI 0 32 6694 2091 1673 53.3%
Marketing MA 15 8 27644 8638 6910 88.4%
Appendix B Results
Table 11: Benchmarking results for the regression tasks. Average mean squared error values over 5 folds and the corresponding standard deviations are reported. Smaller values are better.

### v2

@@ Results
Figure 4 provides a comprehensive ranking of all evaluated methods and their performance over all tasks. The results align with existing literature, highlighting the strong performance of the FT-Transformer architecture (Gorishniy et al.,, 2021), LightGBM, CatBoost and XGBoost (McElfresh et al.,, 2024). CatBoost emerges as the overall best-performing model across all tasks. Among the evaluated models, Mambular stands out as the top-performing neural model on average across all datasets and significantly outperforms XGBoost. Surprisingly, MambAttention, despite combining the second- and fourth-best models, underperforms compared to both. A potential reason is the disruption of information flow, which architectures like Jamba (Lieber et al.,, 2024) counteract by structuring attention layers exclusively after six Mamba blocks and integrating a mixture of experts. However, while this approach enhances expressivity in larger models, it can lead to pronounced overfitting on these comparably smaller tabular datasets.
Detailed results for all datasets and tasks can be found in Table 5 in Appendix A, with additional results on further models provided in Appendix G.
To further validate the potential of autoregressive models as an alternative for tabular deep learning, we compare the top five models on the benchmark suite from Fischer et al., (2023). The results, presented in Figure 5 and in detail in Appendix G in Table 6 , show that Mambular and CatBoost significantly outperform the other models on average. We use the single train-test split provided by Fischer et al., (2023) and exclude datasets already listed in Table 5 to avoid redundancy.
Distributional Regression
To validate Mambular’s suitability for tabular problems, we conducted a small task on distributional regression (Kneib et al.,, 2023). Mambular for Location Scale and Shape (MambularLSS) outperforms XGBoostLSS (März,, 2019) in terms of Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery,, 2007) when minimizing the negative log-likelihood while maintaining a small MSE. A detailed analysis can be found in Appendix D.
@@ 4 Ablation
Model Architecture
This section explores the impact of various elements of Mambular’s architecture, including (i) different pooling techniques, (ii) interaction layers, and (iii) bidirectional processing (Table 1). Transformer networks for natural language processing often use [cls] token embeddings for pooling (Gorishniy et al.,, 2021), a technique that has also proven beneficial in tabular problems (Thielmann et al., 2024b, ). Therefore, this technique is evaluated here. For pooling techniques, we compared Sum-pooling, Max-pooling, Last token pooling – where only the last token in the sequence is passed to the task-specific model head –, and [cls] pooling [^2] Note that [cls] token is appended to the end of each sequence in this implementation. against standard Average-pooling.
Given the significance of feature interactions in tabular problems, we also assessed the effectiveness of a learnable interaction layer between the features. This layer learns an interaction matrix $\mathbf{W}\in\mathbb{R}^{J\times J}$, such that $\text{interactions}=\mathbf{zW}$, where $\mathbf{z}$ is the input feature matrix, before being passed through the SSM. This evaluation was only implemented for the standard Average pooling technique.
Table 3: Analysis of results for CA Housing. Significantly worse results than the default ordering - numerical features: categorical features - and a kernel size of 4, are marked in red. Increasing the kernel size induces positional invariance for features within the sequence.
Model Kernel=4 $\downarrow$ Kernel=J Ordering
[…]
0.177 $\pm$ 0.007 0.160 $\pm$ 0.007 [LO, MA, LA, TR, TB, Po, Hh, MI, OP]
0.175 $\pm$ 0.008 0.173 $\pm$ 0.009 [LO, MA, TR, LA, TB, Po, Hh, MI, OP]
0.194 $\pm$ 0.010 0.169 $\pm$ 0.008 [LO, MA, TR, TB, LA, Po, Hh, MI, OP]
0.196 $\pm$ 0.011 0.161 $\pm$ 0.012 [LO, MA, TR, TB, Po, LA, Hh, MI, OP]
0.194 $\pm$ 0.011 0.173 $\pm$ 0.009 [LO, MA, TR, TB, Po, Hh, LA, MI, OP]
0.195 $\pm$ 0.010 0.169 $\pm$ 0.009 [LO, MA, TR, TB, Po, Hh, MI, LA, OP]
0.194 $\pm$ 0.012 0.172 $\pm$ 0.011 [LO, MA, TR, TB, Po, Hh, MI, OP, LA]
@@
Interestingly, none of the configurations outperformed the basic architecture of average pooling, no interaction, and one-directional processing. Among the pooling strategies, last token pooling and [cls] token pooling performed significantly worse on two out of the four tested datasets. For this ablation study, a 5-fold cross-validation was performed, with the same hyperparameters used across all models. In bi-directional processing, each direction has its own set of learnable parameters, meaning that bi-directional models have additional trainable parameters. All model configurations can be found in Appendix F.
Sequence ordering
Unlike models that leverage attention layers, Mambular is a sequential model. However, tabular data is not inherently sequential – i.e., the order of features in tabular datasets should not matter. Therefore, we examined the significance of variables’ position within the sequence and how their order impacts model performance.
[…]
Model BA $\uparrow$ AD $\uparrow$ AB $\downarrow$ CA $\downarrow$
[…]
@@
In textual data, shuffling the order of words/tokens significantly affects the outcome, and even swapping single words can lead to entirely different contextualized embeddings. Since these contextualized representations are pooled and fed directly to Mambular’s task-specific head, this could also impact performance.
Evaluation experiments were conducted on four real-world datasets and simulated data (see Appendix C). As illustrated in Table 4, we initially confirmed the impact of the kernel size on tabular problems using Mamba’s default kernel size of 4. The findings indicate that the order of sequences does not significantly influence model performance at the 5% level for the selected datasets, even with a relatively small kernel size. However, this is contingent on the data. Strong interaction effects between features that are positioned further apart than the kernel size in the pseudo-sequence can negatively impact model performance, as demonstrated by the results on the California housing dataset.
The positions of the variables Longitude and Latitude appear to directly affect model performance (Table 3). Performance begins to decline significantly when Longitude and Latitude are outside the kernel window. This issue can be entirely resolved by increasing the kernel size to match the sequence length $J$. For a comprehensive analysis, refer to Appendix C.
5 Limitations
The model we have presented has been tested across various datasets and compared against a range of models. However, we have not conducted hyperparameter tuning, as findings from Grinsztajn et al., (2022) and Gorishniy et al., (2021) suggest that most models perform adequately without tuning. These studies indicate that while hyperparameter tuning can enhance performance across all models simultaneously, it does not significantly alter the relative ranking of the models. This suggests that a model that performs best or worst with default configurations will likely retain its ranking even after extensive tuning. Furthermore, McElfresh et al., (2024) reported similar findings, strengthening the notion that hyperparameter tuning benefits most models equally without changing their comparative performance.
The absence of tuning does leave potential for enhancement across all models. However, the default configurations for the comparison models have been extensively tested in numerous studies. It is anticipated that if any model could gain more from hyperparameter tuning, it would be Mambular, due to the lack of extensive literature guiding its default settings. For the comparison models, we made our selections based on literature to ensure default parameters that are meaningful and high-performing. We managed to replicate average results from studies such as Gorishniy et al., (2021) and Grinsztajn et al., (2022). Moreover, key hyperparameters like learning rate, patience, and number of epochs are shared among all models for a more uniform approach. All hyperparameter configurations can be found in Appendix F.
6 Conclusion
We revisited autoregressive models for tabular deep learning and tested the proposed Mambular and MambAttention architectures against common benchmark models. Our work demonstrates the applicability of autoregressive architectures to tabular problems, offering a unique perspective on the interpretation and management of tabular data by treating it as a sequence. Interestingly, our findings suggest that the combination of autoregressive models and attention blocks significantly underperforms their parent models, which rely solely on autoregression or attention. The performance of Mambular, along with its extension MambularLSS, highlights its broad applicability to a wide range of tabular tasks. Additionally, investigating the optimal feature ordering or integrating column-specific information through textual embeddings could further enhance performance. Viewing tabular data as a sequence offers significant benefits for feature-incremental learning: new features can be directly appended to the sequence, eliminating the need to retrain the entire model. While Mamba is still relatively new compared to architectures like the Transformer, its rapid adoption suggests substantial potential for further advancement.
Appendix A Results
Detailed model performances are given below.
Table 5: Benchmarking results for regression and binary classification tasks. Average mean squared error values and area under the curve values over 5 folds and the corresponding standard deviations are reported. Smaller values are better for the regression tasks and larger values are better for the classification tasks. The best performing model is marked in bold. An overview over all datasets for this benchmark is given in the appendix B. The critical difference diagram is shown in Figure 4.
Regression Tasks Classification Tasks
XGBoost 0.019 0.506 0.171 0.528 0.036 0.119 0.024 0.928 0.929 0.845 0.774 0.922
$\pm$ $0$$\pm$ 0.044 $\pm$ 0.007 $\pm$ 0.008 $\pm$ 0.004 $\pm$ 0.024 $\pm$ 0.004 $\pm$ 0.004 $\pm$ 0.002 $\pm$ 0.008 $\pm$ 0.009 $\pm$ 0.002
RF 0.019 0.461 0.183 0.485 0.028 0.121 0.025 0.923 0.896 0.851 0.789 0.917
$\pm$ 0.001 $\pm$ 0.052 $\pm$ 0.008 $\pm$ 0.007 $\pm$ 0.006 $\pm$ 0.018 $\pm$ 0.002 $\pm$ 0.006 $\pm$ 0.002 $\pm$ 0.008 $\pm$ 0.012 $\pm$ 0.004
LightGBM 0.019 0.459 0.171 0.542 0.039 0.112 0.023 0.932 0.929 0.861 0.788 0.927
$\pm$ 0.001 $\pm$ 0.047 $\pm$ 0.007 $\pm$ 0.013 $\pm$ 0.007 $\pm$ 0.018 $\pm$ 0.003 $\pm$ 0.004 $\pm$ 0.001 $\pm$ 0.008 $\pm$ 0.010 $\pm$ 0.001
CatBoost 0.019 0.457 0.169 0.583 0.045 0.106 0.022 0.932 0.927 0.867 0.796 0.926
$\pm$ $0$$\pm$ 0.007 $\pm$ 0.006 $\pm$ 0.006 $\pm$ 0.006 $\pm$ 0.015 $\pm$ 0.001 $\pm$ 0.008 $\pm$ 0.002 $\pm$ 0.006 $\pm$ 0.010 $\pm$ 0.005
FT-Transformer 0.018 0.458 0.169 0.615 0.024 0.111 0.024 0.926 0.926 0.863 0.792 0.916
$\pm$ 0.001 $\pm$ 0.055 $\pm$ 0.006 $\pm$ 0.012 $\pm$ 0.005 $\pm$ 0.014 $\pm$ 0.001 $\pm$ 0.003 $\pm$ 0.002 $\pm$ 0.007 $\pm$ 0.011 $\pm$ 0.003
MLP 0.066 0.462 0.198 0.654 0.764 0.147 0.031 0.895 0.914 0.840 0.793 0.886
$\pm$ 0.003 $\pm$ 0.051 $\pm$ 0.011 $\pm$ 0.013 $\pm$ 0.023 $\pm$ 0.017 $\pm$ 0.001 $\pm$ 0.004 $\pm$ 0.002 $\pm$ 0.005 $\pm$ 0.011 $\pm$ 0.003
ResNet 0.039 0.455 0.178 0.639 0.606 0.141 0.030 0.896 0.917 0.841 0.793 0.889
$\pm$ 0.018 $\pm$ 0.045 $\pm$ 0.006 $\pm$ 0.013 $\pm$ 0.031 $\pm$ 0.017 $\pm$ 0.002 $\pm$ 0.006 $\pm$ 0.002 $\pm$ 0.006 $\pm$ 0.013 $\pm$ 0.003
NODE 0.019 0.431 0.207 0.613 0.045 0.124 0.026 0.914 0.904 0.851 0.790 0.904
$\pm$ $0$$\pm$ 0.052 $\pm$ 0.001 $\pm$ 0.006 $\pm$ 0.007 $\pm$ 0.015 $\pm$ 0.001 $\pm$ 0.008 $\pm$ 0.002 $\pm$ 0.006 $\pm$ 0.010 $\pm$ 0.005
Mambular 0.018 0.452 0.167 0.628 0.035 0.132 0.025 0.927 0.928 0.861 0.796 0.917
$\pm$ $0$$\pm$ 0.043 $\pm$ 0.011 $\pm$ 0.010 $\pm$ 0.005 $\pm$ 0.020 $\pm$ 0.002 $\pm$ 0.006 $\pm$ 0.002 $\pm$ 0.008 $\pm$ 0.013 $\pm$ 0.003
MambAttention 0.018 0.484 0.189 0.638 0.030 0.142 0.026 0.919 0.921 0.857 0.781 0.911
$\pm$ $0$$\pm$ 0.052 $\pm$ 0.006 $\pm$ 0.003 $\pm$ 0.006 $\pm$ 0.024 $\pm$ 0.002 $\pm$ 0.004 $\pm$ 0.002 $\pm$ 0.004 $\pm$ 0.009 $\pm$ 0.001
Further results on a regression benchmark with a single train-test-validation split are reported below. The datasets are taken from Fischer et al., (2023) with datasets already present in the main results excluded.
Table 6: Benchmarking results for the top-5 performing models for regression tasks from (Fischer et al.,, 2023) using a single provided train-test split. The reported values correspond to mean squared error (MSE), where lower values indicate better performance. Mambular and CatBoost achieve the best results, with identical average ranks. The corresponding critical difference diagram is presented in Figure 5.
Model BH ↓ CW ↓ FF ↓ GS ↓ HI ↓ K8 ↓ AV ↓ KC ↓ MH ↓ NP ↓ PP ↓ SA ↓ SG ↓ VT ↓ Rank ↓
Mambular 0.021 0.701 0.272 0.057 0.595 0.168 0.018 0.137 0.085 0.003 0.402 0.015 0.318 0.003 1.79
FT-Transformer 0.028 0.701 0.301 0.205 0.609 0.451 0.089 0.149 0.101 0.009 0.542 0.033 0.360 0.045 4.36
CatBoost 0.032 0.702 0.245 0.041 0.597 0.150 0.004 0.110 0.078 0.005 0.390 0.018 0.297 0.013 1.79
LightGBM 0.048 0.707 0.263 0.059 0.599 0.239 0.024 0.140 0.091 0.009 0.452 0.031 0.302 0.013 3.26
XGBoost 0.039 0.752 0.281 0.078 0.635 0.259 0.004 0.161 0.098 0.006 0.403 0.024 0.329 0.013 3.71
Appendix B Datasets
Table 7: The used datasets for benchmarking. All datasets are taken from the UCI Machine Learning repository. #num and #cat represent the number of numerical and categorical features respectively. The number of features thus determines for Mambular the ”sequence length”. The train, test and validation numbers represent the average number of samples in the respective split for the 5-fold cross validation. Ratio marks the percentage of the dominant class for the binary classification tasks.
Name Abbr. #cat #num train test val ratio
Regression Datasets
Diamonds DI 4 7 34522 10788 8630 -
Abalone AB 1 8 2673 835 668 -
California Housing CA 1 9 13210 4128 3302 -
Wine Quality WI 0 12 4158 1299 1039 -
Parkinsons PA 2 20 3760 1175 940 -
House Sales HS 8 19 13832 4322 3458 -
CPU small CPU 0 13 5243 1638 1310 -
Classification Datasets
Bank BA 13 8 28935 9042 7233 88.3%
Adult AD 9 6 31259 9768 7814 76.1%
Churn CH 3 9 6400 2000 1600 79.6%
FICO FI 0 32 6694 2091 1673 53.3%
Marketing MA 15 8 27644 8638 6910 88.4%
Appendix C Sequence ordering
We test two different shuffling settings: i) shuffling the embeddings after they have passed through the embedding layer, ii) shuffling the sequence of variables before being passed through the embedding layers.
@@ Appendix C Sequence ordering
C.1 California Housing
The p-values for the sequence ordering and positioning of Latitude and Longitude is given below.
Table 9: Detailed analysis of results for CA Housing, including p-statistics.
@@ C.1 California Housing
Given these results, and to verify, that the kernel size of 4 is the cause of this effect, we further analyzed the dataset. Below are more results for Mambular with random shuffling. Again we can see the the position of Latitude and Longitude significantly impact model performance, whenever these two variables are further apart than the fixed kernel size of 4.
Table 10: Analysis of results for CA Housing
Model CA $\downarrow$ $p$-value Ordering
[…]
0.174 $\pm$ 0.009 0.304 [Po, Hh, MI, OP, LO, LA, MA, TR, TB]
0.195 $\pm$ 0.012 0.005 [LO, MA, TR, TB, Po, Hh, MI, OP, LA]
0.197 $\pm$ 0.010 0.002 [MA, LO, TR, TB, Po, Hh, MI, LA, OP]
0.188 $\pm$ 0.010 0.014 [MA, TR, LO, TB, Po, Hh, LA, MI, OP]
0.178 $\pm$ 0.010 0.137 [MA, LO, LA, TR, TB, Po, Hh, MI, OP]
0.177 $\pm$ 0.008 0.142 [MA, TR, TB, Po, LA, LO, Hh, MI, OP]
0.178 $\pm$ 0.009 0.123 [LA, LO, MA, TR, TB, Po, Hh, MI, OP]
0.172 $\pm$ 0.070 0.420 [Hh, TB, Po, MI, MA, OP, LA, LO, TR]
0.177 $\pm$ 0.010 0.171 [LO, Po, OP, LA, MI, MA, TR, Hh, TB]
0.190 $\pm$ 0.010 0.009 [Hh, TB, LO, MI, Po, OP, TR, MA, LA]
Appendix D Distributional Regression
Distributional regression describes regression beyond the mean, i.e., the modeling of all distributional parameters. Thus, Location Scale and Shape (LSS) models can quantify the effects of covariates on not just the mean but also on any parameter of a potentially complex distribution assumed for the responses. A major advantage of these models is their ability to identify changes in all aspects of the response distribution, such as variance, skewness, and tail probabilities, enabling the model to properly disentangling aleatoric uncertainty from epistemic uncertainty.
While this has been a common standard in classical statistical approaches (Stasinopoulos and Rigby,, 2008), it has not yet been widely adopted by the ML community. Recent interpretable approaches (Thielmann et al., 2024a, ), however, have demonstrated the applicability of distributional regression in tabular deep learning. Furthermore, approaches like XGBoostLSS (März,, 2019; März and Kneib,, 2022) demonstrate that tree-based models are capable of effectively solving such tasks. Below, we show that Mambular for Location Scale and Shape (MambularLSS) outperforms XGBoostLSS in terms of Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery,, 2007) when minimizing the negative log-likelihood while maintaining a small MSE.
CRPS
[…]
See Gneiting and Raftery, (2007) for more details.
AB CA
CRPS $\downarrow$ MSE $\downarrow$ CRPS $\downarrow$ MSE $\downarrow$
MambularLSS 0.345 $\pm$ 0.016 0.458 0.201 $\pm$ 0.004 0.181
XGBoostLSS 0.359 $\pm$ 0.016 0.479 0.227 $\pm$ 0.005 0.215
Appendix E MambaTab
In addition to the popular tabular models described above, we tested the architecture proposed by Ahamed and Cheng, 2024a . MambaTab is the first architecture to leverage Mamba blocks for tabular problems. However, the authors propose using a combined linear layer to project all inputs into a single feature representation, transforming the features into a pseudo-sequence of fixed length 1. This approach simplifies the recursive update from Eq. 2 into a matrix multiplication and makes the model resemble a ResNet due to the residual connections in the final processing. Utilizing a sequential model with a sequence length of 1 does not fully exploit the strengths of sequential processing, as it reduces the model’s capacity to capture dependencies across multiple features.
We tested the architecture proposed by Ahamed and Cheng, 2024a and could achieve similar results for shared datasets, but overall found MambaTab to perform similar to a ResNet, aligning with expectations (see Table LABEL:tab:averages and LABEL:tab:reg_regression). Additionally, we experimented with transposing the axes to create an input matrix of shape (1) $\times$ (Batch Size) $\times$ (Embedding Dimension), as outlined in their implementation. While this approach draws on ideas from TabPFN (Hollmann et al.,, 2022), it did not lead to performance improvements in our experiments. When using PLE encodings and increasing the number of layers and dimensions compared to the default implementation from Ahamed and Cheng, 2024a we are able to increase performance.
MambaTab (Ahamed and Cheng, 2024a, ) significantly differs from Mambular, since it is not a sequential model. To achieve the presented results from MambaTab, we have followed the provided implementation from the authors retrieved from https://github.com/Atik-Ahamed/MambaTab. It is worth noting, however, that MambaTab benchmarks the model on a lot of smaller datasets. 50% of the benchmarked datasets have not more than 1000 observations. Additionally, the provided implementation suggests, that MambaTab does indeed not iterate over a pseudo sequence length of 1, but rather over the number of observations, similar to a TabPFN (Hollmann et al.,, 2022). We have also tested this version, denoted as MambaTabT but did not find that it performs better than the described version. On the Adult dataset, our achieved result of 0.901 AUC on average is very similar to the default results reported in Ahamed and Cheng, 2024a with 0.906. The difference could be firstly due to us performing 5-fold cross validation and secondly different seeds in model initialization.